OpenStudy (anonymous):

Question: 50 of 63: Two boxes of apples contain a total of 24 apples. If you halve the number of apples in the first box and add 4 apples to the second box, the total changes to 20 apples. How many apples are in each box initially? Select one of the options below as your answer: A. 12 and 12 B. 18 and 6 C. 16 and 8 D. 10 and 14

OpenStudy (anonymous):

Two boxes of apples contain a total of 24 apples. If you halve the number of apples in the first box and add 4 apples to the second box, the total changes to 20 apples. How many apples are in each box initially? Let number of apple in first box be X. Number of apple in second box = 24 - X 0.5X + (24-x+4) = 20 0.5X - X + 28 = 20 8 = 0.5X X = ???????

OpenStudy (anonymous):

C

OpenStudy (anonymous):

1/2 of 16 is 8 and 8 + 4 is 12 so 8 + 12 = 20